# Feature Attribution with Necessity and Sufficiency via Dual-stage Perturbation Test for Causal Explanation

### v1

@@ Abstract
We investigate the problem of explainability in machine learning. To address this problem, Feature Attribution Methods (FAMs) measure the contribution of each feature through a perturbation test, where the difference in prediction is compared under different perturbations. However, such perturbation tests may not accurately distinguish the contributions of different features, when their change in prediction is the same after perturbation. In order to enhance the ability of FAMs to distinguish different features’ contributions in this challenging setting, we propose to utilize the probability (PNS) that perturbing a feature is a necessary and sufficient cause for the prediction to change as a measure of feature importance. Our approach, Feature Attribution with Necessity and Sufficiency (FANS), computes the PNS via a perturbation test involving two stages (factual and interventional). In practice, to generate counterfactual samples, we use a resampling-based approach on the observed samples to approximate the required conditional distribution. Finally, we combine FANS and gradient-based optimization to extract the subset with the largest PNS. We demonstrate that FANS outperforms existing feature attribution methods on six benchmarks.
@@ 1 Introduction
Feature attribution is a method for explaining machine learning (ML) models by assigning weights to input features, where the absolute value of these weights denotes their contribution to a model’s prediction. Mathematically, suppose we have a function $f$ that denotes a trained ML model, a target input $\mathbf{x}$$=$$(x_{1},\cdots,x_{d})$ to be explained, and a baseline $\mathbf{x}^{\prime}$, where $x^{\prime}_{i}$ is the approximate value that feature $x_{i}$ would take if it were considered missing. An attribution of the prediction at input $\mathbf{x}$ relative to $\mathbf{x}^{\prime}$ is a vector $\alpha_{f}(\mathbf{x},\mathbf{x}^{\prime})$$=$$(a_{1},...,a_{d})$, where $a_{i}$ is the contribution of $x_{i}$ to the prediction $f(\mathbf{x})$.
@@
Standard feature attribution methods (FAMs) measure the contribution of each feature through a perturbation test, i.e., comparing the difference in prediction under different perturbations, or before and after the perturbation, where a subset of the variables of $\mathbf{X}$ are set to their baseline values (Lundberg & Lee, 2017; Sundararajan et al., 2017). However, this perturbation test may not accurately distinguish the contribution of different features to the prediction when their change in prediction is the same after perturbation. To better understand this phenomenon, we provide an example as follows.
Example 1.1.
@@
Consider a binary classification model $f$ for $\mathbf{X}$$=(X_{1},X_{2},X_{3})$in $\mathbb{R}^{3}$ that is captured by a function $Y$=$h(X_{1},X_{2},X_{3})$=$\mathbb{I}(X_{1}$--$X_{2}$$>$1)1)1 ), where $Y$ is a binary variable representing true or false. Assuming that the baseline $\mathbf{x}^{\prime}$ is $(0,0,0)$, given the target input $\mathbf{x}=(1,1,1)$.
@@
The attribution scores obtained from the existing FAMs for features $x_{1}$ and $x_{3}$ may both be 00 since the model’s predictions given $\mathbf{x}$ under different perturbations relative to $\mathbf{x}^{\prime}$ for any subset of $\mathbf{X}$ are always equal to $h(\mathbf{x})$, which may misleadingly suggest that $x_{1},x_{3}$ contribute equally to $h(\mathbf{x})$. Clearly, the contributions of $x_{1},x_{3}$ are not equal, since the value of $h(\mathbf{x})$ is determined solely by the difference between $x_{1}$ and $x_{2}$ being less than $1$, and is not influenced by $x_{3}$. To solve this problem, we propose to use the Probability of Necessity and Sufficiency (PNS). Let $\mathbf{s}$ denote a subset of dimensions of $\mathbf{X}$. We measure the importance of the feature subsets of $\mathbf{x}$ on $\mathbf{s}$ (denoted as $\mathbf{x}_{\mathbf{s}}$, e.g., $\mathbf{x}_{\mathbf{s}}$=$\{x_{1},x_{3}\}$ for $\mathbf{s}$=$\{1,3\}$) by measuring the probability that perturbing $\mathbf{x}_{\mathbf{s}}$ is a necessary and sufficient cause for the prediction to change. Specifically, given any two events $A_{\mathbf{s}}$ and $B_{\mathbf{s}}$, PNS measures the probability that event $\bar{B}$ occurs when event $\bar{A}$ occurs, and event B occurs when event A occurs (as the “if and only if” of the relationship). In our context, we take event $A$ as perturbing the features on $\mathbf{s}$ of different samples, which share the features $\mathbf{x}_{\mathbf{s}}$ and event $B$ as prediction changes after perturbation. In Example 1.1, the PNS for $x_{3}$ (i.e. $\mathbf{s}=\{3\}$) is 0, since perturbing $x_{3}$ is impossible to cause the prediction to change, while the PNS for $x_{1}$ may be greater than 0 because, for any sample that shares the value of $x_{1}$, the prediction may change or remain unchanged when $x_{1}$ is perturbed or not perturbed.
Based on the above analysis, to enhance the ability of the FAMs to distinguish different feature contributions when the prediction changes are the same after perturbation, we present a Feature Attribution with Necessity and Sufficiency (FANS) method to measure the importance of each feature subset. First, to estimate PNS, we design a general perturbative prediction process of ML models accounting for input, prediction, and perturbation. Second, we calculate the PNS for each feature subset as its importance score. To achieve this goal, we first decompose PNS into a weighted sum of the probability that event $A$ is a sufficient cause for event $B$ (PS) and the probability that event $A$ is a necessary cause for event $B$ (PN). Then we design two dual-stage perturbation tests for PN and PS respectively. Specifically, to estimate PS, in the factual stage, we draw inputs from a distribution conditional on the fact that predictions remain unchanged after perturbing the features of $\mathbf{X}$ on $\bar{\mathbf{S}}$, where $\mathbf{X}_{\mathbf{S}}$=$\mathbf{x}_{\mathbf{S}}$. In the interventional stage, we perturb the collected inputs on $\mathbf{S}$ and count the proportion of prediction changes. Similarly, the perturbation test for PN is constructed by exchanging the perturbation and prediction events involved in the above two stages . Since the conditional distribution in the factual stage may be complex, we propose to use Sampling-Importance-Resampling on the observed inputs to approximate it. Further, we combine FANS and a gradient-based optimization method to extract the subset with the largest PNS. We demonstrate FANS outperforms existing FAMs on six benchmarks of image and graph data.
2 Related Work
2.1 ML Explainability
Explainable artificial intelligence (XAI) methods aim to provide human-readable explanations to help users comprehend and trust the outputs created by ML models. At present, a series of XAI methods have been developed, mainly including building interpretable models (e.g., linear regression model) and generating post hoc explanations. The post hoc explanation methods mainly include feature attribution methods (Yang et al., 2023b), which provide an estimate of the contribution of input features to a model’s prediction, example-based methods (Koh & Liang, 2017), which provide example(s) that may be considered similar or as prototype(s), hidden semantics (Simonyan et al., 2013), which make sense of certain hidden neurons, etc. In this paper, we focus on the feature attribution approach.
Existing feature attribution methods (FAMs) mainly measure the contribution of features to prediction through a perturbation test, i.e., by perturbing the features and comparing the resulting differences in predictions. Typically, Shapley value-based methods (Shapley et al., 1953; Lundberg & Lee, 2017) measure the difference in a model’s predictions given data points with and without the feature included. Gradient-based methods (Sundararajan et al., 2017; Yang et al., 2023b) measure the rate of change in prediction corresponding to a tiny change in an input feature. LIME-based methods (Ribeiro et al., 2016; Dhurandhar et al., 2022) learn simplified, interpretable models on perturbed feature subsets. There are also many methods to find the feature subset with the highest contribution by optimizing a loss function between the $L_{p}$ norm of an input feature mask and the prediction given the masked inputs (Tan et al., 2022; Fong & Vedaldi, 2017). However, current feature attribution methods may not accurately distinguish the contributions between different features when their change in prediction is the same after perturbation. To address this issue, FANS employs a dual-stage perturbation test, which has advantages in handling these cases.
2.2 Probability of Necessity and Sufficiency
PNS (Pearl, 2022) can be used to measure the “if and only if” of the relationship between two events and currently has many applications. For example, in ML explainability, LEWIS (Galhotra et al., 2021), LENS (Watson et al., 2021) and NSEG (Cai et al., 2022) use sufficiency or necessity to measure the contribution of input feature subsets. However, LEWIS uses only PS or PN, while LENS uses a variant of PS, so the attribution scores of these methods lack sufficient and necessary theoretical guarantees. NSEG extracts the feature subset with the largest PNS by introducing an identifiability assumption and then optimizing the lower bound of PNS, but our FANS does not require this assumption and directly uses PNS to measure the importance of each feature subset. In the out-of-distribution generalization problem, it can be used to extract domain-invariant information (Yang et al., 2023a). In the causal effect estimation problem, it can be used to learn individual responses from population data (Mueller et al., 2021).
3 Feature Attribution with Model Perturbative Prediction Process
Our goal is to quantify the impact of features on the resultant predictions by fixing certain features to values similar to those observed in a target input $\mathbf{x}$, and then allowing other features to change according to the data distribution. Figure 1 illustrates this process. Specifically, let $\mathbf{X}$ denote the input variable and $\mathbf{S}$ denote a subset of dimensions on which to perturb the input, with $\bar{\mathbf{S}}$ denoting its complement. Sequentially, given a constant $b\geq 0$, define the distribution of $\mathbf{X}$ with the condition that the $L_{p}$ norm of the value of $\mathbf{X}$ on $\mathbf{S}$ with respect to the value of the target input $\mathbf{x}$ is not greater than $b$. To avoid confusion, we represent the corresponding random variables that follow this conditional distribution as $\tilde{\mathbf{X}}$:
$\small\tilde{\mathbf{X}}\sim P(\mathbf{X}~{}|~{}\|\mathbf{X}_{\mathbf{S}}-\mathbf{x}_{\mathbf{S}}\|_{p}\leq b).$ (1)
Note that the conditional distribution in Eq. 1 can be obtained by sampling from the original input data distribution given a value of S and retaining only the samples that satisfy the condition $|\mathbf{X}_{\mathbf{S}}-\mathbf{x}_{\mathbf{S}}|_{p}\leq b$. Finally, given the ML model $f$, a baseline $\mathbf{x}^{\prime}$, a $d$-dimensional Bernoulli distribution with hyperparameter $\bm{\phi}$ specified by the user (it is fixed to $0.5$ in practice), let $\bm{1}$ denote an all-ones vector and we design a perturbation function $g$ to randomly perturb the features of $\tilde{\mathbf{x}}$ specified by $\mathbf{S}$ as follows.
@@ 3 Feature Attribution with Model Perturbative Prediction Process
where $\circ$ represents element-wise multiplication. The reason for the Bernoulli distribution introduced in Eq. 2 is that in the process of quantifying the model’s predictions by the impact of perturbing the dimensions specified by $\mathbf{s}$, the dimensions that can influence the model predictions may be $\mathbf{s}$ or a subset of $\mathbf{s}$. Therefore, we sample selected dimensions of $\mathbf{s}$ to obtain a more comprehensive quantification. Further, we use $\mathbf{Y}$=$f(g(\tilde{\mathbf{x}},\mathbf{S},\mathbf{x}^{\prime}))$ to denote the process of perturbing the values of $\tilde{\mathbf{x}}$ and then predicting.
Remark 3.1.
Our model perturbative prediction process differs from the standard causal generation process in two ways. First, variable $\mathbf{Y}$ is the model’s prediction rather than a property of the real-world data. Second, by intervening on a variable in our process, the outcome of its corresponding children’s variables is available.
Figure 1: Model perturbative prediction process. $\mathbf{S}$ denotes a subset of dimensions of $\mathbf{X}$ for perturbation. $\tilde{\mathbf{X}}$ represents an input with fixed features on $\mathbf{S}$ that are similar to the target input $\mathbf{x}$.
Given any model $f$, the goal of feature attribution can be formalized with the help of the aforementioned perturbative prediction process. Specifically, given a $d$-dimensional vector $\mathbf{x}$, we let $z(\mathbf{x},\mathbf{s})$ be a perturbative prediction process, so the goal of feature attribution is to devise a score function $\phi:z,x_{\mathbf{s}}\rightarrow w_{\mathbf{s}}$, such that this score function $\phi$ is able to distinguish features with significantly different contributions to model predictions.
@@ 4 Feature Attribution as a Problem of PNS Measurement
Based on the previous problem definition, we find that the goal of feature attribution is to find a score function $g$ that can distinguish features with significantly different contributions to model predictions. However, it is not a trivial task since the degree of change in predictions and the contribution of features to predictions may be asymmetric.
To solve this problem, one straightforward intuition is to describe the features that influence the prediction if and only if they are present, which can be addressed with the help of the probability of necessity and sufficiency. Therefore, we reformalize PNS as follows. Suppose $f$ is a binary classification model. Standard feature attribution methods measure the contribution of each feature to the prediction by comparing different perturbations of the feature $\mathbf{x}$, which involves two events $A_{\mathbf{s}},B_{\mathbf{s}}$. Specifically, we let $A_{\mathbf{s}}$ denote an event of perturbation on the feature subset $\mathbf{s}~{}of~{}\mathbf{X}$ relative to $\mathbf{x}^{\prime}$ and $B_{\mathbf{s}}$ denote an event of prediction change on the prediction $y$ with $|y^{\prime}-y|>c$ and $y^{\prime}$ is the value of $y$ after some perturbation on the subspace ${\mathbf{s}}$ relative to $\mathbf{x}^{\prime}$. Overloading the notation, we let $A_{\mathbf{s}}$ be the feature with the index of ${\mathbf{s}}$ that is perturbed, and $B_{\mathbf{s}}$ be the label that is changed after some perturbations. Moreover, we have the following definitions (Pearl, 2022).
@@
$PN=P(\bar{B_{\mathbf{s}}}_{\bar{A}_{\mathbf{s}}}|A_{\mathbf{s}},B_{\mathbf{s}}),$ (3)
PN is the probability that, given that events $A_{\mathbf{s}}$ and $B_{\mathbf{s}}$ both occur initially, event $B_{\mathbf{s}}$ does not occur after event $A_{\mathbf{s}}$ is changed from occurring to not occurring.
@@
$PS=P({B_{\mathbf{s}}}_{A_{\mathbf{s}}}|\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})$ (4)
PS is the probability that, given that events $A_{\mathbf{s}}$ and $B_{\mathbf{s}}$ both did not occur initially, event $B_{\mathbf{s}}$ occur after event $A$ is changed from not occurring to occurring.
@@
$PNS=PN\cdot P(A_{\mathbf{s}},B_{\mathbf{s}})+PS\cdot P(\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})$ (5)
@@
PNS is the sum of PN and PS, each multiplied by the probability of its corresponding condition. PNS measures the probability that event $A_{\mathbf{s}}$ is a necessary and sufficient cause for event $B_{\mathbf{s}}$.
When evaluating the feature weight $w_{\mathbf{s}}$ in a certain input, if we hope that $w_{\mathbf{s}}$ can reflect the probability that perturbing $x_{\mathbf{s}}$ is a sufficient and necessary cause of prediction change, then this $w_{\mathbf{s}}$ can be defined as:
$\small w_{\mathbf{s}}=P(\bar{B}_{{\mathbf{s}}_{\bar{A}_{\mathbf{s}}}}|A_{\mathbf{s}},B_{\mathbf{s}})\cdot P(A_{\mathbf{s}},B_{\mathbf{s}})+P(B_{{\mathbf{s}}_{A_{\mathbf{s}}}}|\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})\cdot P(\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}}).$ (6)

### v3

@@ Abstract
We investigate the problem of explainability for machine learning models, focusing on Feature Attribution Methods (FAMs) that evaluate feature importance through perturbation tests. Despite their utility, FAMs struggle to distinguish the contributions of different features, when their prediction changes are similar after perturbation. To enhance FAMs’ discriminative power, we introduce Feature Attribution with Necessity and Sufficiency (FANS), which find a neighborhood of the input such that perturbing samples within this neighborhood have a high Probability of being Necessity and Sufficiency (PNS) cause for the change in predictions, and use this PNS as the importance of the feature. Specifically, FANS compute this PNS via a heuristic strategy for estimating the neighborhood and a perturbation test involving two stages (factual and interventional) for counterfactual reasoning. To generate counterfactual samples, we use a resampling-based approach on the observed samples to approximate the required conditional distribution. We demonstrate that FANS outperforms existing attribution methods on six benchmarks. Please refer to the source code via https://github.com/DMIRLAB-Group/FANS.
@@ 1 Introduction
Feature attribution is a method for explaining machine learning (ML) models by assigning weights to input features, where the absolute value of these weights denotes their contribution to a model’s prediction. Mathematically, suppose we have a function $f$ that denotes a trained ML model, a target input $\mathbf{x}^{t}$$=$$(x^{t}_{1},\cdots,x^{t}_{d})$ to be explained, and a baseline $\mathbf{x}^{\prime}$$=$$(x^{\prime}_{1},\cdots,x^{\prime}_{d})$ , where $x^{\prime}_{i}$ is the approximate value that feature $x^{t}_{i}$ would take if it were considered missing. An attribution of the prediction at input $\mathbf{x}^{t}$ relative to $\mathbf{x}^{\prime}$ is a vector $\alpha_{f}(\mathbf{x}^{t},\mathbf{x}^{\prime})$$=$$(a_{1},...,a_{d})$, where $a_{i}$ is the contribution of $x^{t}_{i}$ to the prediction $f(\mathbf{x}^{t})$.
@@
Standard feature attribution methods (FAMs) measure the contribution of each feature through a perturbation test, i.e., comparing the difference in prediction under different perturbations, e.g., Shapley Values (Shapley et al., 1953), or before and after the perturbation, e.g., gradient or surrogate-based methods (Lundberg & Lee, 2017; Sundararajan et al., 2017), where a subset of the variables of $\mathbf{X}$ are set to their baseline values.
However, this perturbation test may not accurately distinguish the contribution of different features to the prediction when their changes in prediction are similar after perturbation. To better understand this phenomenon, we provide an example as follows.
@@
Example 1. Consider a binary classification model $f$ for $\mathbf{X}$$=(X_{1},X_{2},X_{3})$in $\mathbb{R}^{3}$ that is captured by a function $Y$=$h(X_{1},X_{2},X_{3})$=$\mathbb{I}(X_{1}$--$X_{2}$$>$1)1)1 ), where $Y$ is a binary variable representing true or false. Assume the baseline $\mathbf{x}^{\prime}$$=$ $(0,0,0)$ and the target input $\mathbf{x}^{t}$$=$(1,1,1)111(1,1,1)( 1 , 1 , 1 ).
@@
The attribution scores for features $x^{t}_{1}$ and $x^{t}_{3}$ obtained from the existing FAMs may both be 00, as the model’s predictions given $\mathbf{x}^{t}$ under different perturbations relative to $\mathbf{x}^{\prime}$ are always equal to $h(\mathbf{x}^{t})$, which may misleadingly suggest that $x^{t}_{1},x^{t}_{3}$ contribute equally to $h(\mathbf{x}^{t})$. Clearly, the contributions of $x^{t}_{1},x^{t}_{3}$ are not equal, since the value of $h(\mathbf{x}^{t})$ is determined solely by the difference between $x^{t}_{1}$ and $x^{t}_{2}$ being less than $1$, and is not influenced by $x^{t}_{3}$. In other words, the causes for the predictions for $x^{t}_{1}$ and $x^{t}_{3}$ remaining unchanged after perturbation are different.
Figure 1: Causal diagram of standard perturbation test in feature attribution. $\mathbf{S}$ denotes a subset of dimensions of $\mathbf{X}$ for perturbation. $\tilde{\mathbf{X}}$ represents an input with fixed features on $\mathbf{S}$ that are similar to the target input $\mathbf{x}^{t}$.
Based on this key observation, we extend feature attribution by finding a neighborhood of the input $\mathbf{x}^{t}$ containing $x^{t}_{3}$ (or $x^{t}_{2}$) such that perturbing samples within this neighborhood has the highest probability of causation called Probability of Necessity and Sufficiency (PNS) of two counterfactual questions about $x^{t}_{3}$, and we take this PNS as the importance of $x^{t}_{3}$. The counterfactual questions are as follows. “for the inputs within this neighborhood, and the predictions for these inputs remain unchanged when some features (except $x^{t}_{3}$) are perturbed, would the predictions change if $x^{t}_{3}$ were perturbed?” and “for the inputs within other neighborhoods containing $x^{t}_{3}$, and the predictions for these inputs change when only $x^{t}_{3}$ is perturbed, would the predictions remain unchanged if some features (except $x^{t}_{3}$) were perturbed?” The first question assesses the sufficiency of perturbing $x^{t}_{3}$ (cause A) for the prediction change (effect B), i.e., the ability of cause A to produce effect B. Similarly, the second question assesses the necessity, i.e., the dependence of effect B on cause A (Pearl, 2000). In Example 1, the PNS for $x^{t}_{3}$ is 0, since perturbing $x^{t}_{3}$ is impossible to cause the prediction to change, while the highest PNS for $x^{t}_{1}$ may be greater than 0 because the prediction may change when $x^{t}_{1}$ is perturbed.
The above analysis demonstrates that evaluating the maximum PNS for prediction change more effectively distinguishes the importance of features than the extent of those changes. Hence, we develop a novel Feature Attribution with Necessity and Sufficiency (FANS) to compute the maximum PNS of the above two counterfactual questions about each feature subset. Specifically, we abstract the common understanding of the perturbation test in feature attribution by proposing a structural causal model (SCM). We then propose a heuristic strategy to calculate a neighborhood of $\mathbf{x}^{t}$ that is used to estimate the highest PNS of the counterfactual questions for each feature subset. To quantify this PNS, we then propose a novel dual-stage perturbation test to implement the counterfactual reasoning paradigm Abduction-Action-Prediction (Pearl, 2013). The first stage is to generate samples from a distribution condition on the factual perturbation and prediction of the counterfactual question (i.e., Abduction). The second stage involves perturbing the feature subset of each sample that differs from their factual perturbation (i.e., Action) and then updating the predictions. Finally, we infer the PNS based on the proportion of changes and remaining unchanged in the prediction. However, the conditional distribution in the first stage may be complex, we thus propose to use Sampling-Importance-Resampling (DB, 1988) on the observed inputs to approximate it. Further, we combine FANS and a gradient-based optimization method to extract the subset with the largest PNS. We demonstrate FANS outperforms existing FAMs on six benchmarks of image and graph data.
2 Causal Model for Feature Attribution
In this section, we develop a principled framework to model the perturbation test in feature attribution, and describe the motivation of our proposed FANS approach that will be formally introduced in Section 3. The essential notations are presented in Table 3 of the Appendix.
Structural Causal Model (SCM) for standard perturbation test. In a SCM, all dependencies are generated by functions that compute variables from other variables. In this paper, we take SCM as our starting point and try to develop everything from there. Figure 1 illustrates the causal diagram of the standard perturbation test in feature attribution. Let $\mathbf{X}$ denote the $d$-dimensional input variable and $\mathbf{S}$ denotes a subset of dimensions $\{1,2,...,d\}$ of the input, with $\bar{\mathbf{S}}$ denoting its complement. We first observe the causal relationships $\mathbf{S}\to\tilde{\mathbf{X}}$ and $\mathbf{X}\to\tilde{\mathbf{X}}$. In this case, we sample around the target input $\mathbf{x}^{t}$ with the aim of taking into account the local behavior of the black box model (Dhurandhar et al., 2022). Mathematically, let $b$ denote the local boundary around $\mathbf{x}^{t}$ and we define the neighborhood of $\mathbf{x}^{t}$ as the distribution of $\mathbf{X}$ with the condition that the $L_{p}$ norm of $\mathbf{X}$ on $\mathbf{S}$ with respect to the value of $\mathbf{x}^{t}$ is not greater than $b$. To avoid confusion, we represent the corresponding variable as $\tilde{\mathbf{X}}$, as follows:
$\small\tilde{\mathbf{X}}\sim P(\mathbf{X}~{}|~{}\|\mathbf{X}_{\mathbf{S}}-\mathbf{x}^{t}_{\mathbf{S}}\|_{p}\leq b).$ (1)
Thus, the question is how to approximate the local boundary $b$ accurately. Our main idea is to define the boundary $b$ that perturbing samples within that boundary has a high probability of being Necessity and Sufficiency cause for the change in predictions. The formal definition of boundary $b$ will be introduced in Definition 3.4. Note that the distribution in Eq. 1 can be empirically approximated by sampling from the distribution of input given a value of $\mathbf{S}$ and selecting samples that meet the condition $\|\mathbf{X}_{\mathbf{S}}-\mathbf{x}^{t}_{\mathbf{S}}\|_{p}\leq b$. We then observe the causal relationships $\tilde{\mathbf{X}}\to\mathbf{Y}$ and $\mathbf{S}\to\mathbf{Y}$. In this case, we perturb the features of $\tilde{\mathbf{X}}$ on $\mathbf{S}$ by a perturbation function $g$, and then feed the perturbed $\tilde{\mathbf{X}}$ into the model to generate a new prediction $\mathbf{Y}$, which is denoted by
$\mathbf{Y}=f(g(\tilde{\mathbf{X}},\mathbf{S},\mathbf{x}^{\prime})),$ (2)
where we choose the most commonly used settings for baseline $\mathbf{x}^{\prime}$ based on the data type and these details can be found in Appendix C.4. Given a $d$-dimensional Bernoulli distribution with a hyperparameter $\varphi$ (fixed to $0.5$ in practice) and an all-one vector $\bm{1}$, we implement the function $g$ to randomly perturb the features of $\tilde{\mathbf{X}}$ on $\mathbf{S}$ as follows.
@@ 2 Causal Model for Feature Attribution
where $\circ$ represents element-wise multiplication. The reason for the Bernoulli distribution introduced in Eq. 3 is that in the process of quantifying the model’s predictions by the impact of perturbing the dimensions specified by $\mathbf{S}$, the dimensions that can influence the model predictions may be $\mathbf{S}$ or a subset of $\mathbf{S}$. Therefore, we sample selected dimensions of $\mathbf{S}$ to obtain a more comprehensive quantification.
Goal of standard perturbation test. A FAM using a standard perturbation test aims to quantify the importance of a feature subset corresponding to $\mathbf{s}$ (denoted by $a_{\mathbf{s}}$) which comprises two steps: (1) evaluate two model predictions of $\mathbf{Y}$ under different $\mathbf{S}$ values ($\mathbf{s}$ and $\mathbf{s}^{\prime}$) through the causal model mentioned above, where $\mathbf{s}^{\prime}$ used for reference is generally an empty set (i.e., no perturbation); (2) compute the importance $a_{\mathbf{s}}$ by comparing the difference in predictions. Formally, $a_{\mathbf{s}}$ is given by a function $\phi:\mathcal{Y}\times\mathcal{Y}\to\mathcal{A}$ that maps two predictions in $\mathcal{Y}$ to a scalar in $\mathcal{A}\subseteq\mathbb{R}$.
Remark 1. Most of the perturbation tests designed by various FAMs are compatible with our introduced causal model, contingent on their respective implementations of $\tilde{\mathbf{X}}$, score function $\phi$ and $\mathbf{s}^{\prime}$. For example, Shapley Values (Shapley et al., 1953) fixes the variable $\tilde{\mathbf{X}}$ to be the target input $\mathbf{x}^{t}$, defines $\phi$ as Shapley Value, and sets $\mathbf{s}=\mathbf{s}^{\prime}\cup x^{t}_{i}$.
Remark 2. Our causal model differs from the standard causal generation process in two ways. First, variable $\mathbf{Y}$ is the model’s prediction rather than a property of the real-world data. Second, by intervening on a variable in our process, the outcome of its corresponding children’s variables is available.
@@ 3 Feature Attribution as a Problem of PNS Measurement
[…]
@@
$PN=P(\bar{B}_{{\mathbf{s},c}_{\bar{A}_{\mathbf{s},b}}}|A_{\mathbf{s},b},B_{\mathbf{s},c}),$ (4)
PN represents the probability of the following counterfactual question: given that events $A_{\mathbf{s},b}$ and $B_{\mathbf{s},c}$ both occur initially, would $B_{\mathbf{s},c}$ not occur if $A_{\mathbf{s},b}$ were changed from occurring to not occurring?
@@
$PS=P({B_{\mathbf{s},c}}_{A_{\mathbf{s},b}}|\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})$ (5)
PS represents the probability of the following counterfactual question: given that events $A_{\mathbf{s},b}$ and $B_{\mathbf{s},c}$ both did not occur initially, would $B_{\mathbf{s},c}$ occur if $A_{\mathbf{s},b}$ were changed from not occurring to occurring?
@@
$PNS=PN\cdot P(A_{\mathbf{s},b},B_{\mathbf{s},c})+PS\cdot P(\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c}).$ (6)
@@
PNS is the weighted sum of PN and PS, each multiplied by the probability of its corresponding condition. PNS measures the probability that $A_{\mathbf{s},b}$ is a necessary and sufficient cause for $B_{\mathbf{s},c}$ (as the “if and only if” of the relationship).
When evaluating the importance $w_{\mathbf{s}}$ of the input $\mathbf{x}^{t}$ on the dimension subset $\mathbf{s}$, we want $w_{\mathbf{s}}$ to reflect the probability of causation, then this objective can be formalized as follows: find a neighborhood for the target input $\mathbf{x}^{t}$ such that, compared to other neighborhoods of $\mathbf{x}^{t}$, perturbing samples within this neighborhood on $\mathbf{s}$ have the highest probability of being necessary and sufficient cause for the prediction change, and we take this probability as $w_{\mathbf{s}}$. Mathematically, the objective is to find the boundary $b$ and the threshold $c$ for events $A_{\mathbf{s},b}$ and $B_{\mathbf{s},c}$ that maximize the PNS, as illustrated in Definition 3.4, with an illustrative example provided.
Definition 3.4.
(Necessary and Sufficient Attribution) Necessary and Sufficient Attribution of the input $\mathbf{x}^{t}$ on the dimension subset $\mathbf{s}$ is defined as
$\small w_{\mathbf{s}}:=\max_{b,c}PN\cdot P(A_{\mathbf{s},b},B_{\mathbf{s},c})+PS\cdot P(\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c}).$ (7)
Example 2. Consider a $1$-dimensional model, $y$$=$$\sigma(\mathbf{x}^{2})$, with a target sample $\mathbf{x}^{t}$$=$$0.2$, a baseline $\mathbf{x}^{\prime}$$=$00, $\mathbf{s}$$=$$\{1\}$, $b$$=$$0.01$, and $c$$=$$0.1$. Since $\bar{\mathbf{s}}$ is empty, the domain of definition for $P(\tilde{\mathbf{X}}|\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})$ is empty. Thus, we focus on calculating PN and $P(A_{\mathbf{s},b},B_{\mathbf{s},c})$. Consider no perturbation will be performed when calculating PN since $\bar{\mathbf{s}}$ is empty. Consequently, PN equals 1. $P(A_{\mathbf{s},b},B_{\mathbf{s},c})$ can be estimated through the empirical distribution of the dataset. Overall, the importance score of $\mathbf{x}^{t}$ is equal to $P(A_{\mathbf{s},b},B_{\mathbf{s},c})$.